How much of a change in GDP will result if government increases its spending by $3 billion and the MPC in the economy is 0.67?


1
Expert's answer
2021-08-15T17:06:02-0400

"Change in GDP=\frac{government spending}{1-MPC}"


"=\frac{ \$3 billion}{1-0.67}"


"=\frac{\$3billion}{0.33}"


"=\$9.09"

Hence increase in GDP is $9.09 billion